# Global identification of dynamic panel models with interactive effects

### v1

@@ 4.3 Proof of Theorem 1
Recall that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)=0$ must hold in order for Case 3 to be plausible. There are two ways in which $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)$ can be 0: either all the coefficients are equal to 0, so that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial; or the coefficients of the polynomial are such that real roots exist and at least one of them coincide with $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$.
We can see that Assumption 6 implies that $a=b=c=0$ can only occur on a set with probability 0, as it involves the factor realizations lying on a particular low-dimensional manifold (see Lemma 2). That is, the first case has zero probability of occurring.
We next show that the second case also has zero probability of occurring. To see this, note that $\widetilde{J}_{(2,3),(1,4)}$ is a polynomial of order 2 in $\widetilde{\alpha}$. Its roots are:
@@
Examining the expressions of $a$, $b$, and $c$ and making use of Lemma 2 and Assumption 6, we conclude that these events can never occur with positive probability (the factors $f_{2},f_{3},f_{4}$ must lie exactly in a lower dimensional manifold). Hence $\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$ is only possible on a set of zero probability.
@@
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$
Our goal is to show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$ can only occur with probability zero, implying that $\widetilde{\alpha}=\alpha$ must hold almost surely. Recall from Proposition 2 that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree in between $1$ and $12\bar{r}+$ 1, given by:
@@
By Assumptions 5 and 6, $\alpha,\mathbf{D}\ \text{and}\ \mathbf{\Psi}$ are deterministic population parameters, and $\mathbf{F}$ is drawn from a continuous distribution. The condition $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$ can occur in two possible ways: either all of the coefficients of this random polynomial are equal to 0, or by setting $\widetilde{\alpha}$ equal to one of the roots of this polynomial $\widetilde{J}_{R,C}(\widetilde{\alpha}^{\star},\theta)=0$ (by the fundamental theorem of algebra we know that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ has at most $12\bar{r}+$ 1 real roots).
Consider the first case. Let $\kappa\in\{1,2,\dots,2\bar{r}+1\}$ be the degree of the random polynomial $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$. In order for all the coefficients to be equal to 0, the factor realizations must be such that the coefficients $a_{0}(\theta),a_{1}(\theta),\dots,a_{\kappa}(\theta)$ are all simultaneously equal to 0. By Lemma 2, this is only possible in a set of probability 0.
@@
Now consider the second case. If $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ has no real roots, then it is clear that it must be that $\widetilde{\alpha}=\alpha$ in order for (15) to hold, so the proof would be over. Now, assume $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ has at least one real root. Pick any of these roots, say $\widetilde{\alpha}^{\star}_{R,C}(\theta)$. By definition, $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ solves:
$\widetilde{J}_{R,C}\left(\widetilde{\alpha}^{\star}_{R,C},(\theta),\theta\right)=0$
@@
Note that $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ is a function of the other true data-generating process parameters and can be obtained from the above condition. Moreover, it also depends on the rows $R$ and columns $C$ chosen as in general the random polynomials $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ are different (except for a set of factor realizations of probability zero).
@@
As we used the diagonal exclusion minor associated with rows and columns $R,C$ to pin down $\widetilde{\alpha}$ we have no other free parameter, so this same $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ must satisfy:
$\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}(\theta),\theta)=0$
@@
for any other valid set of $R^{\prime},C^{\prime}$. This can only occur in two different ways: either $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial, or the random polynomial $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ shares at least one root with $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$.
@@
As we have previously discussed, the case in which $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial is an event that occurs with probability zero. Therefore, consider the second case. We know that the coefficients of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ can be exactly equal only in a set of probability zero, so the polynomials are almost surely different. Then, the case in which they share the root $\widetilde{\alpha}^{\star}_{R,C}(\theta)$ can be expressed as:
@@
$\begin{split}\widetilde{J}_{R^{\prime},C^{\prime}}&=\sum_{j=1}^{\bar{r}+1}(-1)^{\bar{r}+1+j^{{}^{\prime}}}\Bigg{[}J_{r_{\bar{r}+1}^{{}^{\prime}},c_{j}^{{}^{\prime}}}\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}}_{R^{\prime}-r_{\bar{r}+1}^{{}^{\prime}},C^{\prime}-c_{j}^{{}^{\prime}}}\right)+\Omega_{r_{\bar{r}+1}^{{}^{\prime}},c_{j}^{{}^{\prime}}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{{}^{\prime}},C^{\prime}-c_{j}^{{}^{\prime}}}\\ &\quad+\left(\alpha-\widetilde{\alpha}^{\star}_{R,C}(\theta)\right)J_{r_{\bar{r}+1}^{{}^{\prime}},c_{j}^{{}^{\prime}}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{{}^{\prime}},C^{\prime}-c_{j}^{{}^{\prime}}}\Bigg{]}=0\end{split}$ (18)
@@
Note that the above condition depends only on the structural parameters of the model. Moreover, it defines a particular low-dimensional manifold in which the random factors $\mathbf{F}$ must lie in order for both polynomials to share the root $\widetilde{\alpha}^{\star}_{R,C}(\theta)$. However, by Assumption 6 and Lemma 2, the event in which the random factors lie in this particular low-dimensional manifold occurs with probability zero.
@@
Then, the case $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$ can only occur in a set of probability 0. Thus, it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to be satisfied. This implies that $\alpha$ is almost surely globally identified. $\square$
4.4 Global identification of $\mathbf{F}$, $\mathbf{\Psi}$ and $\mathbf{D}$
@@
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\Psi\mathbf{F^{\prime}}$, where $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ must be a $T\times T$ symmetric PSD matrix of rank $\bar{r}$ and $\mathbf{\widetilde{D}}$ a $T\times T$ PD diagonal matrix.
Under the Assumptions of our model, this is the case shown by Anderson and Rubin (1956), so all the rest of the parameters are almost surely globally identified as well. $\square$
@@ 5 Individual fixed-effects
So far, we have established that the model parameters are globally identified under the assumption that all factors in $\mathbf{F}$ follow a continuous probability distribution (Assumption 6). While this assumption is quite general and encompasses many relevant applications, it notably excludes one important case: individual fixed-effects.
$y_{it}=\alpha y_{it-1}+\delta_{t}+\gamma_{i}+\lambda_{i}^{{}^{\prime}}f_{t}+\varepsilon_{it}$ (19)
where $\gamma_{i}$ denotes the individual fixed effect for observation $i$, and the rest of the parameters are defined as in (1). It is important to note that we will not estimate the fixed effects $\gamma_{i}$, doing so will cause the incidental parameters problem. We simply treat $\gamma_{i}$ as the factor loadings of a special factor, which taking on the value of 1 for all $t$. The number of factors in this model is $\bar{r}=dim(f_{t})+1$.
Using the same procedure as in Section 2.1, we project the first observation $y_{i1}$ onto the vector $[1,\gamma_{i},\lambda_{i}]$, yielding the following expression: $y_{i1}=\delta_{1}^{*}+\gamma_{i}f_{\gamma}^{*}+\lambda_{i}^{\prime}f_{1}^{*}+\varepsilon_{i1}^{*}$, where $(\delta_{1}^{*},f_{\gamma}^{*},f_{1}^{*})$ are the projection coefficients and $\varepsilon_{i1}^{*}$ is the projection residual. For notational simplicity, we drop the asterisks from the coefficients and from the residual, as these are free parameters, and we do not require the error terms $\varepsilon_{it}$ to have identical distributions across $t$. Therefore, for $t=1$, we can rewrite the equation as:
@@
Combining this with equation (19) for $t\geq 2$ and stacking the observations over $t$, we obtain:
@@
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}^{{}^{\prime}}\\ 1&f_{2}^{{}^{\prime}}\\ \vdots&\vdots\\ 1&f_{T}^{{}^{\prime}}\end{bmatrix},\quad\text{ and }\eta_{i}=\begin{bmatrix}\gamma_{i}\\ \lambda_{i}\end{bmatrix}$
Note that the first element in the first column of $\mathbf{F}$ is $f_{\gamma}$, not 1. Importantly, the first column of $\mathbf{F}$ consists of many 1’s, which are fixed constants and follow degenerate distributions. As a result, the factor matrix does not satisfy the conditions of Assumption 6. However, since this assumption is a sufficient but not necessary condition, we will show that the model remains globally identified. The key insight is that, as long as there is sufficient variation in some of the factors, identification can still be achieved.
@@
Note that the presence of individual fixed-effects imposes structure on $\mathbf{F}$, so we require less restrictions to eliminate rotational indeterminacy. We only need to impose two restrictions on $\mathbf{F}$ to prevent indeterminacy (see Appendix B for detailed explanation). We shall impose the normalization in the following way, referred to as tail normalization:
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}\\ 1&f_{2}\\ \vdots&\vdots\\ 1&f_{T-2}\\ 1&0\\ 1&1\end{bmatrix}$
This normalization restricts the last two realizations of the factor $f$ to be 0 and 1 respectively. This is purely for convenience in term of exposition. Normalizing other entries (except the first row) has the same effect. Under this normalization, $\mathbf{F}A=\mathbf{F}$, if and only if $A=I_{2}$, thus eliminating rotational indeterminacy.
Before showing that $\alpha$ can be almost surely identified in this context we shall explicitly state the identification assumptions:
Assumption 1’.
@@
Assumption 2’.
@@
Assumption 3’.
$\eta_{i}$ are either random or fixed constants with sample covariance matrix $\bm{\Psi}_{N}$ satisfying $\bm{\Psi}_{N}\to\bm{\Psi}$ with $\bm{\Psi}$ being positive definite, as $N\to\infty$.
Assumption 4’.
$\alpha,\mathbf{\Psi},d_{1},d_{2},\dots,d_{T}$ are fixed population parameters.
Assumption 5’.
@@
Assumption 6’.
$d_{t}\neq\alpha d_{t-1}$ when $\alpha\neq 1$.
Assumptions 1’-5’ are an adaptation of the ones outlined in Section 3 for the particular case we are studying. While we conjecture that $f_{\gamma}$ drawn from a continuous distribution is not strictly needed to attain identification, it substantially simplifies the proof exposition.
Assumption 6’ is imposed because our strategy is not able to show identification in that particular case (it will become clearer why during the proof). We argue that this does not undermine the generality of our results for two main reasons. First, the case in which $d_{t}=\alpha d_{t-1}$ with $\alpha\neq 1$ is a restriction on the values of the parameters that have 0-lebesgue measure in the parameter space. Second, and more importantly, it is hard to think any true data generating process such that the variance of the idiosyncratic errors should evolve as $d_{t}=\alpha d_{t-1}$ when $\alpha\neq 1$, so this is a case we believe is never observed in practice. Note that the case in which $\alpha=1$ and $d_{t}=\alpha d_{t-1}$ is identified. This implies that if our process exhibits an unit root and homoskedastic variance we can almost surely identify $\alpha$, which is a case that should have more empirical relevance.
Under this assumptions, we can show that $\alpha$ is almost surely globally identified:
@@
Let $y_{it}$ be generated from (19). Then, under Assumptions 1’-6’, $\alpha$ is almost surely globally identified.
Proof: The proof follows in a similar way as in the previous section, as most of the results derived did not depend on the distributional assumption imposed on $\mathbf{F}$. For instance, the definition of $\mathbf{O}$ is the same as before (for the case of $\bar{r}=2$) and Claim 1 still holds, so for any $3\times 3$ diagonal exclusion minor of $\mathbf{O}$ we should have:
@@
$\det\left(\mathbf{M}^{\mathbf{O},3}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$
We will show that $\alpha$ is almost surely identified when $T=6$. This is a sufficient condition for identification for any bigger $T$, as it is clear that having a higher number of observations can only help identification, not make it worse.
Let $T=6$ and $R=(1,2,3)$, $C=(4,5,6)$. The determinant of the diagonal exclusion minor associated with these rows and columns is given by:
$\det\left(\mathbf{M}^{\mathbf{O},3}_{(1,2,3),(4,5,6)}\right)$ $=(\alpha-\widetilde{\alpha})\left(\Psi_{1,1}\Psi_{2,2}-\Psi_{1,2}^{2}\right)$ (20)
×$\quad\times\left(d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma}\right)$
×$\quad\times(\widetilde{\alpha}f_{4}+1)(\widetilde{\alpha}-1)=0$
where, in this case, $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of order 2. The proof is similar to the case of $\bar{r}=1$ in Section 4.3. Note that for equation (20) to hold, one of the following five alternatives must be true:

### v2

@@ 4.3 Proof of Theorem 1
Recall that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)=0$ must hold in order for Case 3 to be plausible. There are two ways in which $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ can be 0: either all the coefficients are equal to 0, so that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial; or the coefficients of the polynomial are such that real roots exist and at least one of them coincide with $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$.
We can see that Assumption A7 implies that $a=b=c=0$ can only occur on a set with probability 0, as it involves the factor realizations lying on a particular low-dimensional manifold (see Lemma 2). That is, the first case has zero probability of occurring.
We next show that the second case also has zero probability of occurring. To see this, note that $\widetilde{J}_{(2,3),(1,4)}(\widetilde{\alpha},\theta^{0})$ is a polynomial of degree 2 in $\widetilde{\alpha}$. Its roots are:
@@
Examining the expressions of $a$, $b$, and $c$ and making use of Lemma 2 and Assumption A7, we conclude that these events can never occur with positive probability (the factors $f_{2},f_{3},f_{4}$ must lie exactly in a lower dimensional manifold). Hence $\widetilde{\alpha}=\frac{f_{4}}{f_{3}}$ is only possible on a set of zero probability.
@@
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Our goal is to show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$ for all possible valid combinations of $R$ and $C$ can only occur with probability zero, implying that $\widetilde{\alpha}=\alpha$ must hold almost surely. Recall from Proposition 2 that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree in between $1$ and $2\bar{r}+1$, given by:
@@
By Assumptions A6 and A7, $\alpha,\mathbf{D}\ \text{and}\ \mathbf{\Psi}$ are deterministic population parameters, and $\mathbf{F}$ is drawn from a continuous distribution. The condition $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$ can occur in two possible ways: either all of the coefficients of this random polynomial are equal to 0, or by setting $\widetilde{\alpha}$ equal to one of the roots of this polynomial $\widetilde{J}_{R,C}\left(\widetilde{\alpha}^{\star},\theta^{0}\right)=0$ (by the fundamental theorem of algebra we know that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ has at most $2\bar{r}+1$ real roots).
The first case is automatically ruled-out by Proposition 2, as it states that $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is a random polynomial in $\widetilde{\alpha}$ of degree in between $1$ and $2\bar{r}+1$ almost surely. This result guarantees that with probability one $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is not the zero polynomial.
@@
Now consider the second case. If $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ has no real roots, then it is clear that it must be that $\widetilde{\alpha}=\alpha$ in order for (15) to hold, so the proof would be over. Now, assume $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ has at least one real root. Pick any of these roots, say $\widetilde{\alpha}^{\star}_{R,C}(\theta^{0})$. By definition, $\widetilde{\alpha}^{\star}_{R,C}(\theta^{0})$ solves:
$\widetilde{J}_{R,C}\left(\widetilde{\alpha}^{\star}_{R,C}(\theta^{0}),\theta^{0}\right)=0$
@@
Note that $\widetilde{\alpha}^{\star}_{R,C}(\theta^{0})$ is a function of the other true data-generating process parameters and can be obtained from the above condition. Moreover, it also depends on the rows $R$ and columns $C$ chosen as in general the random polynomials $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ are different (except for a set of factor realizations of probability zero).
@@
As we used the diagonal exclusion minor associated with rows and columns $R,C$ to pin down $\widetilde{\alpha}$ we have no other free parameter, so this same $\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)$ must satisfy:
$\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})=0$
@@
for any other valid set of $R^{\prime},C^{\prime}$. This can only occur in two different ways: either $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial, or the random polynomial $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ shares at least one root with $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$.
@@
As we have previously discussed, the case in which $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial is an event that occurs with probability zero. Therefore, consider the second case. We know that the coefficients of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ and $\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ can be exactly equal only in a set of probability zero, so the polynomials are almost surely different. Then, the case in which they share the root $\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)$ can be expressed as:
@@
$\begin{split}\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})&=\sum_{j=1}^{\bar{r}+1}(-1)^{\bar{r}+1+j}\Bigg[J_{r_{\bar{r}+1}^{\prime},c_{j}^{\prime}}\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}}_{R^{\prime}-r_{\bar{r}+1}^{\prime},C^{\prime}-c_{j}^{\prime}}\right)\\ &+\quad\Omega_{r_{\bar{r}+1}^{\prime},c_{j}^{\prime}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{\prime},C^{\prime}-c_{j}^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})\\ &+\quad\left(\alpha-\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)\right)J_{r_{\bar{r}+1}^{\prime},c_{j}^{\prime}}\widetilde{J}_{R^{\prime}-r_{\bar{r}+1}^{\prime},C^{\prime}-c_{j}^{\prime}}(\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right),\theta^{0})\Bigg]=0\end{split}$ (18)
@@
Note that the above condition depends only on the structural parameters of the model. Moreover, it defines a particular low-dimensional manifold in which the random factors $\mathbf{F}$ must lie in order for both polynomials to share the root $\widetilde{\alpha}^{\star}_{R,C}\left(\theta^{0}\right)$. However, by Assumption A7 and Lemma 2, the event in which the random factors lie in this particular low-dimensional manifold occurs with probability zero.
@@
Then, the case $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$ can only occur in a set of probability 0. Thus, it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to be satisfied. This implies that $\alpha$ is almost surely globally identified. $\square$
4.4 Global identification of $\mathbf{F}$, $\mathbf{\Psi}$, and $\mathbf{D}$
@@
if and only if $\mathbf{\widetilde{D}}=\mathbf{D}$, and $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$, where $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ must be a $T\times T$ symmetric PSD matrix of rank $\bar{r}$ and $\mathbf{\widetilde{D}}$ a $T\times T$ PD diagonal matrix.
Under the Assumptions of our model, this is the case shown by Anderson and Rubin, (1956), so all the rest of the parameters are almost surely globally identified as well. $\square$
@@ 5 Individual fixed-effects
Consider a panel data model with time, individual and interactive effects:
$y_{it}=\alpha y_{it-1}+\delta_{t}+\gamma_{i}+\lambda_{i}^{\prime}f_{t}+\varepsilon_{it}$ (19)
where $\gamma_{i}$ denotes the individual fixed effect for observation $i$, and the rest of the parameters are defined as in (1). It is a well-known result that individual fixed-effects can be absorbed into the factor structure by treating $\gamma_{i}$ as the factor loadings of a special factor, which takes the value of 1 for all $t$. As a result, the factor matrix does not satisfy the conditions of Assumption A7 because the factor associated with the individual fixed-effects follows a degenerate distribution.
Therefore, the results derived in Theorem 1 do not apply in this case. We can overcome this issue in two ways. The first approach consists of showing that our proof can be adapted to allow for one factor with a degenerate distribution. In practice, global identification requires that there is enough variation in some of the factors, but there could be some other factors following degenerate distributions. It is important to note that in this case we do not estimate $\gamma_{i}$ (doing so would cause incidental parameter problems), but we absorb it in the factor structure and identify the variance of $\gamma_{i}$ and the covariance with the rest of the factor loadings through $\mathbf{\Psi}$.
The second approach consists of differencing the data to get rid of the individual fixed effects as in Hayakawa et al., (2023). While Hayakawa et al., (2023) claims that global identification of $\alpha$ cannot be attained in the differenced data model we show that their conjecture is incorrect and that we can use a slightly modified version of our proof to derive global identification when we work in differences.
In practice, we advocate for working with the data in levels. Differencing the data eliminates important information about the process contained when the model is studied in levels. Moreover, it introduces time dependence between the errors as a mechanical result from differencing the idiosyncratic error. It also requires additional periods for identification as a result of both losing one observation by differencing and the temporal dependence between the errors.
5.1 Analysis in levels
We start from the model (19) and absorb the individual fixed effects into the factor structure:
$y_{it}=\alpha y_{it-1}+\delta_{t}+\eta_{i}^{\prime}F_{t}+\varepsilon_{it}$ (20)
where everything is defined as in (1) but $\eta_{i}=[\gamma_{i},\lambda_{i}]$ and $F_{t}=[1,f_{t}]$. Using the same procedure as in Section 2.1, we project the first observation $y_{i1}$ onto the vector $1,\gamma_{i},\lambda_{i}]$, yielding the following expression: $y_{i1}=\delta_{1}^{*}+\gamma_{i}f_{\gamma}^{*}+\lambda_{i}^{\prime}f_{1}^{*}+\varepsilon_{i1}^{*}$, where $(\delta_{1}^{*},f_{\gamma}^{*},f_{1}^{*})$ are the projection coefficients and $\varepsilon_{i1}^{*}$ is the projection residual. For notational simplicity, we drop the asterisks from the coefficients and from the residual, as these are free parameters, and we do not require the error terms $\varepsilon_{it}$ to have identical distributions across $t$. Therefore, for $t=1$, we can rewrite the equation as:
@@ 5.1 Analysis in levels
Combining this with equation (20) and stacking the observations over $t$, we obtain:
@@
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}^{\prime}\\ 1&f_{2}^{\prime}\\ \vdots&\vdots\\ 1&f_{T}^{\prime}\end{bmatrix},\quad\text{ and }\eta_{i}=\begin{bmatrix}\gamma_{i}\\ \lambda_{i}\end{bmatrix}$
Note that the first element in the first column of $\mathbf{F}$ is $f_{\gamma}$, not 1. Importantly, the first column of $\mathbf{F}$ consists of many 1’s, which are fixed constants and follow degenerate distributions. As a result, the factor matrix does not satisfy the conditions of Assumption A7. However, since this assumption is a sufficient but not necessary condition, we show that the model remains globally identified. The key insight is that, as long as there is sufficient variation in some of the factors, identification can still be achieved.
@@
Note that the presence of individual fixed-effects imposes structure on $\mathbf{F}$, so we require fewer restrictions to eliminate rotational indeterminacy. We only need to impose two restrictions on $\mathbf{F}$ to prevent indeterminacy (see Appendix B for detailed explanation). We shall impose the normalization in the following way, referred to as tail normalization:
$\mathbf{F}=\begin{bmatrix}f_{\gamma}&f_{1}\\ 1&f_{2}\\ \vdots&\vdots\\ 1&f_{T-2}\\ 1&0\\ 1&1\end{bmatrix}$ (21)
This normalization restricts the last two realizations of the factor $f$ to be 0 and 1 respectively. This is purely for convenience in terms of exposition. Normalizing other entries (except the first row) has the same effect. Under this normalization, $\mathbf{F}\mathbf{A}=\mathbf{F}$, if and only if $\mathbf{A}=\mathbf{I}_{2}$, thus eliminating rotational indeterminacy.
Before showing that $\alpha$ can be almost surely identified in this context, we shall explicitly state the identification assumptions:
Assumption B1.
@@
Assumption B2.
$\mathbf{\Psi}$ is an unrestricted symmetric and PD matrix, and $\mathbf{F}$ is normalized as (21)
Assumption B3.
@@
Assumption B4.
$\varepsilon_{i}$ are i.i.d. over $i$. $\mathbb{E}[\varepsilon_{it}]=0$, $\mathbb{V}(\varepsilon_{it})=d_{t}>0$, and $\varepsilon_{it}$ are independent over $t$. $\varepsilon_{it}$ is independent from the factor loadings $\eta_{i}$ and the factors $\mathbf{F}$.
Assumption B5.
$\eta_{i}$ are either random or fixed constants with sample covariance matrix $\mathbf{\Psi}_{N}$ satisfying $\mathbf{\Psi}_{N}\to\mathbf{\Psi}$ with $\mathbf{\Psi}$ being PD, as $N\to\infty$.
Assumption B6.
$\alpha,d_{1},d_{2},\dots,d_{T},\mathbf{\Psi}$ are fixed population parameters.
Assumption B7.
@@
Assumption B8.
There exists a time period $t$ such that $d_{t}\neq\alpha d_{t-1}$ when $\alpha\neq 1$.
Assumptions B1-B7 are an adaptation of the ones outlined in Section 3 for the particular case we are studying. While we conjecture that $f_{\gamma}$ drawn from a continuous distribution is not strictly needed to attain identification, it substantially simplifies the proof exposition.
Assumption B8 is imposed because our strategy is not able to show identification in that particular case (it will become clearer why during the proof). We argue that this does not undermine the generality of our results for two main reasons. First, the case in which $d_{t}=\alpha d_{t-1}$ with $\alpha\neq 1$ is a restriction on the values of the parameters that have Lebesgue measure zero in the parameter space. Second, and more importantly, it is hard to think any true data generating process such that the variance of the idiosyncratic errors should evolve as $d_{t}=\alpha d_{t-1}$ when $\alpha\neq 1$, so this is a case unlikely to be of practical relevance. Note that the case in which $\alpha=1$ and $d_{t}=d_{t-1}$ is identifiable. Thus, when for $\alpha=1$, the model is identifiable under both homoskedasticity and heterokedasticity.
Under these assumptions, we can show that $\alpha$ is almost surely globally identified:
@@
Let $y_{it}$ be generated from (19). Then, under Assumptions B1-B8, $\alpha$ is almost surely globally identified.
Proof.
The proof follows in a similar way as in the previous section, as most of the results derived did not depend on the distributional assumption imposed on $\mathbf{F}$. For instance, the definition of $\mathbf{O}$ is the same as before (for the case of $\bar{r}=2$) and Claim 1 still holds, so for any $3\times 3$ diagonal exclusion minor of $\mathbf{O}$ we should have:
@@
$\det\left(\mathbf{M}^{\mathbf{O},3}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
We show that $\alpha$ is almost surely identified when $T=6$. This is a sufficient condition for identification for any bigger $T$, since having more observations can only help identification, not make it worse. Let $T=6$ and $R=(1,2,3)$, $C=(4,5,6)$. The determinant of the diagonal exclusion minor associated with these rows and columns is given by:
[…]
$\quad\times\left(d_{3}f_{1}-\alpha d_{2}f_{1}-d_{3}f_{2}f_{\gamma}+\alpha^{2}d_{1}f_{2}-\alpha^{2}d_{1}f_{3}+\alpha d_{2}f_{3}f_{\gamma}\right)$
$\quad\times(\widetilde{\alpha}f_{4}+1)(\widetilde{\alpha}-1)=0$
where, in this case, $\widetilde{J}_{(1,2,3),(4,5,6)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree 2. The proof is similar to the case of $\bar{r}=1$ in Section 4.3. Note that for (22) to hold, one of the following five alternatives must be true: